OpenStudy (anonymous):

J is the midpoint of . The coordinates of K are (-8, 7) and the coordinates of L are (3, 4), find the coordinates of J (leave in fraction form).

OpenStudy (anonymous):

Find the average of the x-coordinates and the average of the y coordinates. That is your answer.

OpenStudy (anonymous):

Remember to find an average of two numbers you add them together and divide by 2.

OpenStudy (anonymous):

so \[\frac{ -8+3 }{ 2 }=\frac{ -5 }{ 2 }\] and \[\frac{ 7+4 }{ 2 }=\frac{ 11 }{ 2 }\] So the midpoint is \[J\left( \frac{ -5 }{ 2 }, \frac{ 11 }{ 2 } \right)\]